You need a capacitance of 55.0 ?F, but you don't happen to have a 55.0 ?F capacitor. You do have a 85.0 ?F capacitor.

Part A

What additional capacitor do you need to produce a total capacitance of 55.0 ?F?

Express your answer with the appropriate units.

Part B

Should you join the two capacitors in parallel or in series?
Answer #1

We have a capacitor of capacitance = 85 F (make unit as per question as it is showing as ? here)

needed capacitance = 55 F

additional capacitor needed is to be added in series as the total capacitance needed less than available capacitance

so

1/Creq = 1/Cgiven + 1/C

1/55 = 1/85 + 1/C

1/C = 1/55 - 1/85

C = 155.833 F

additional capacitor do we need = 155.833 F